OpenStudy (gabylovesyou):

What are the x-intercept(s) of the graph of y + 30 = x^2 + x?

OpenStudy (gabylovesyou):

@HawkCrimson

OpenStudy (anonymous):

it's 2x ??

OpenStudy (gabylovesyou):

x^2

OpenStudy (gabylovesyou):

@moli1993

OpenStudy (anonymous):

0,-30

OpenStudy (gabylovesyou):

These are my choices: (−5, 0) and (6, 0) (−5, 0) and (−6, 0) (5, 0) and (−6, 0) (5, 0) and (6, 0)

OpenStudy (gabylovesyou):

@moli1993

OpenStudy (raden):

to get the x-intercept(s), take y =0

OpenStudy (anonymous):

oops . . i was wrng took x=0 instead :p

OpenStudy (gabylovesyou):

@RadEn i got C.. :)

OpenStudy (raden):

30 = x^2 + x subtract by 30 to both sides : 30-30 = x^2+x-30 x^2+x-30 = 0 factor out! (x+6)(x-5) =0 for the zeroes, take each factor be equal zero x+6=0, we have a solution x=-6 x-5=0, the solution is x=5 so, they are (-6,0)and (5,0)

OpenStudy (raden):

yeah, it is C